# SIR-RL: Reinforcement Learning for Optimized Policy Control during Epidemiological Outbreaks in Emerging Market and Developing Economies

### v1

@@ 1 Introduction
The need for a nuanced understanding of how interventions impact both health outcomes and economic indicators became increasingly evident, prompting a comprehensive examination by epidemiologists to assist policymakers [7]. The outbreak of COVID-19 has prompted epidemiologists to research on various aspects, including mobility control [8, 9], vaccination strategies [10, 11], non-pharmaceutical interventions (NPIs) like restricting population movements and gatherings, closing schools and businesses, and requiring masks indoors [12, 13, 14], and financial considerations [15]. Despite the numerous studies conducted, very few explore how common interventions meet multiple policy objectives or how a precise articulation of the main policy goals directs the selection of the most effective interventions in terms of health and economic results [16, 8, 17, 18, 19, 20, 21, 22]. The economic impact of the COVID-19 pandemic varied between rich and poor countries. Although COVID-19 deaths had a slightly larger negative effect on the Gross Domestic Product (GDP) in advanced economies, this difference was not statistically significant. However, lockdown restrictions were found to have a more damaging impact on economic activity in emerging and developing economies [6, 23, 24]. It’s also suggested that an increase in COVID-19 cases was associated with the introduction of harsher NPIs and lockdown measures could be relaxed once vaccination rates increase [23, 25].
@@
Many economists have studied the effect of COVID-19 on the economy of nations [6, 26, 27, 28]. In advanced economies like Korea, where the stringency index was below the median the recession was milder than other advanced economies like the United Kingdom where the stringency was much higher [26], they achieved it mostly with very aggressive testing, contact tracing, and enforced quarantines [29, 30]. In India, social distancing and containment measures have been effective in reducing the number of COVID-19 cases but have come with economic costs. Social distancing had the most adverse effect on the economy in areas with high urbanization [27].
@@
In this paper we optimize the government policies regarding stringency as it controls both the spread of the disease and the economy. To model the epidemiological data [31] we use the simple SIR model without vital dynamics [32, 33, 34] as it is assumed that the timescale is small enough that it can be neglected [35]. By lesioning the model, as opposed to proposing a new mathematical model with more specialized compartments to more accurately represent the actual environment [36, 37], we effectively model the disease progression. Our model (SIR with lockdown and time-varying vaccination rate) builds on the foundational SIR model, by accounting for the recovery reached through vaccination [38, 39, 40, 41, 42] and the effects of lockdown [43, 44, 21, 45]. Although the traditional SIR model is a valuable tool for understanding the spread of infectious diseases, it assumes that parameters like the transmission rate ($\beta$) and recovery rate ($\gamma$) are constant over time, which may not always be the case. In this paper, we propose a more sophisticated approach by introducing a time-dependent SIR model [46], enabling us to account for the changing dynamics of the pandemic due to factors such as lockdowns and vaccination rates. This proposed model effectively addresses the real-world conditions acts as a solution that is both effective and extendable. However, the study has limitations; First, the deterministic SIR model (predecessor to our proposed model) fails to account for chance in disease spread and lacks confidence intervals on results and while stochastic models incorporate chance, they are typically more challenging to analyze than their deterministic counterparts [33]. Secondly, the underreporting of cases during the period selected by our study. Lastly, the reinforcement learning agent should be resistant to how the vaccination rate changes, and different values for $\beta$ and $\gamma$ – keeping them the same scopes the environment for succumbing to wishful thinking which can be potentially dangerous. Therefore, before an actual deployment of the model, it would be a good measure to introduce stochasticity to these parameters ($\beta$ and $\gamma$) and the vaccination rate ($\nu$).
@@
In summary, by using reinforcement learning augmented with deep learning techniques for the SIR with lockdown and time-varying vaccination rate environment, we can better understand the effects of lockdown measures on both public health outcomes and economic stability during a pandemic. However, it is crucial to consider the limitations of this approach and take into account a comprehensive set of decision factors in order to make informed policy decisions that are tailored to specific socioeconomic contexts.
@@ 2.1 Simple SIR Model
During an epidemic, the fundamental drivers of an epidemic growth is the rate of infection $\beta$ i.e., the average number of infections per infected case and the infectious period $1/\gamma$ i.e., the average period for which the infected case is infected for. Epidemics can only happen if the case is infectious enough for long enough and this defined by $R_{0}=\beta/\gamma$. Here, $R_{0}$ is the average number of secondary infections caused by each infected case, in an otherwise fully susceptible population.
At the peak of an epidemic there is a decline as there are no more susceptible people left in the pool, therefore, $R_{e}$ (effective reproductive number) comes into play. $R_{e}$ is defined as the average number of secondary cases arising from an infected case, at a given point in an epidemic, therefore, it takes into account the existing immunity of the system [62].
@@
$S$ is the number of susceptible people, $N$ is the total population. At the start of an epidemic when everyone is susceptible, $R_{e}=R_{0}$ as, $S=N$ (i.e., the whole population is susceptible). $\beta$ and $\gamma$ are also used to define probability of and infectious individual infecting another individual $\beta/(\beta+\gamma)$ and the probability of recovery, $\gamma/(\beta+\gamma)$.
Most government policies look at the value of $R_{e}$ to come us with an effective strategy to combat the disease as the fate of the evolution of the disease depends upon it. When $R_{e}$ is less than one, the infected population $I$ will steadily decline to zero. Conversely, if $R_{e}$ is greater than one, the infected population will increase. In other words, when $\frac{dI(t)}{dt}<0\Rightarrow R_{e}<1$ and $\frac{dI(t)}{dt}>0\Rightarrow R_{e}>1$, therefore, the effective reproductive rate $R_{e}$ serves as a critical threshold that determines whether an infectious disease will rapidly extinguish or escalate into an epidemic [35].
To estimate the parameters $\beta$ and $\gamma$ for India, based on the data from May, 2020, to October, 2022, we simply define two cost functions eqs. 8 and 9 to calibrate the model with the use of huber loss [63]. Our model typically uses eq. 8, considering all three compartments: susceptible, infected, and recovered. However, there are instances where we need to balance modeling all groups and focusing on the infected group (population that drives disease spread). In such cases, we consider losses from both eqs. 8 and 9. A weighted sum of the loss functions allows for trade-offs between comprehensive modeling and focusing on infected group dynamics (see section 2.4, where a hyperparameter (window length) is selected keeping both these losses in mind).
@@
Where, $S$ is the number of susceptible people and $\hat{S}$ is the predicted number of susceptible people, similarly, $I$ for infected and $\hat{I}$ for the predicted number of infected people, and $R$ for recovered. Using the equations Equations 13, we minimize the cost function eq. 13 using the Nelder-Mead method [64] to estimate the parameters like $\beta$, $\gamma$, and fit the model to actual data. The following parameters and loss is obtained:
@@ 2.2 SIR Model with Lockdown
Now that, a simple SIR model has been established – we need to model the effects of the stringency index (measure for the strictness of lockdown) on $\beta$ (the effective transmission rate). To do this, we say the flow of susceptibles not only depend on $\beta$ but also $s(t)$ the stringency index at time [45, 65, 44, 66, 22]. The stringency index is a composite measure based on nine response indicators including school closures, workplace closures, and travel bans, rescaled to a value from 0 to 100 (100 = strictest) [67]. This index simply records the strictness of government policies and does not measure or imply the appropriateness or effectiveness of a country’s response i.e., a higher score does not necessarily mean that a country’s response is “better” than others lower on the index.
@@ 2.4 Optimizing Window Length for Time-varying Vaccination Rate
However, as observed by the value of $\nu_{optimal}$ from eq. 28 which is almost negligible and the overestimation of infected individuals in fig. 6(b) suggests that $\nu$ might be varying with time. This suggests to accurately estimate the infected population a time-varying vaccination rate should be used as the transition from susceptibility to direct recovery fluctuates with time [24, 38]. We first find the optimal window length [68] for which the value of $\nu$ is constant for a time sub-interval and results in the least loss from eqs. 8 and 9. For this we use different window lengths ($window\_lengths=5,10,15\dots 40,45,50\textrm{ days}$).
@@
Using the model described by Equations 3234 for each $window\_length_{i}$ in $window\_lengths$, where $i=1,2,3\dots 10$, we calculate the $time\_varying\_\nu$. For each $time\_varying\_\nu$ we then calculate the $loss\_SIR$ and $loss\_I$. This is done as follows:
@@
Figure 1: Loss for Different Window Lengths. We try different window lengths to find the optimal loss for both cases, either when predicting all three populations(susceptible, infected, recovered) or just the infected population.
The results in fig. 1, indicate that a window length of $10$ days yields the least overall loss for all three population groups. However, this window length results in a poor approximation for just the infected group, which is crucial for accurately modelling the spread of the disease. Consequently, we have decided to use a window length of $15$ days, which provides a more accurate approximation for the infected population while still maintaining reasonable loss for the other groups.
@@
This fig. 2 shows us that the $\nu$ coincides with the actual of data of when the vaccination drive was first launched in India [69]. Therefore, using these values we finally, recompute $\beta_{optimal}$ and $\gamma_{optimal}$ by supplying them into the equations for the SIR Model with lockdown and time-varying $\nu$.
@@ 2.6 Modelling Normalized GDP with Stringency
Now, that a relation between $\beta$ and $s(t)$ is set up, it must be investigated how stringency index affects the normalized gross domestic product (GDP) [70, 71]. To do this a polynomial equation of the third degree is fitted to the data points $f(x)=ax^{3}+bx^{2}+cx+d=y$, here, $x$ is the stringency ($s$), and $y$ the normalized GDP, and we minimize the squared error to find the values of coefficients $a,b,c,d$. For India after fitting a 3 degree polynomial, the following equation is obtained:
@@ 2.7 Reinforcement Learning
  • Actions $\altmathcal{A}$: The stringency index variable was analyzed with a sample size of 915. The mean value was approximately $61.96505$, with a standard deviation of $17.66983$. The minimum value was $31.48$, while the maximum value reached $96.3$. And the differences between two consecutive stringencies had a mean of $-0.070919$, and standard deviation of $1.42715$, with the minimum being $-14.36$ and maximum $16.67$. Based on this we define the discrete action space. There are 7 actions for the agent, it can keep the stringency index same, reduce/increase by 2.5, reduce/increase by 5, and reduce/increase by 10 given that the stringency index doesn’t exceed 100 or go below 0.
@@ 2.7.1 Defining the Reward Function
The stringency index emerges as a critical factor influencing both the normalized GDP and the rate of infection spread. The decision to escalate or de-escalate the stringency index is a strategic one, with significant implications. Increasing the stringency decreases the spread of the infection. Conversely, it must be noted that herd immunity can only be achieved when the epidemic reaches its peak i.e., when the effective reproductive number is equal to one ($R_{e}=1$). This can only happen by lowering the stringency index which would allow the natural dynamics of the epidemic to transpire such that the population of susceptible individuals has depleted enough such that it is insufficient to propagate the disease further. Therefore, stringency is used to control the number of infected people and slow down the rate at which the epidemic reaches its peak, so that hospitals could house the number of infected people.
In reinforcement learning, positive rewards promote and negative rewards demote actions. The agent tries to generate such a policy/knowledge to avoid the discouraging situation by following the policy. By designing a proper reward function, it is possible to generate such an agent that may follow the human desired situation. While designing a reward function it is important to note that the rewards we set up truly indicate what we want accomplished. In particular, the reward signal is not the place to impart to the agent prior knowledge about how to achieve what we want it to do [52]. Taking inspiration from similar work [19], we define the reward function.
@@ 2.7.2 Deep Reinforcement Learning and Training
The agent observes the percentage of the population that is susceptible, infected, recovered, and time-varying data like the GDP, and previous actions taken to change the stringency, and $R_{e}$. Since Stable Baselines3 can support multiple inputs (time-series data, single data points and images) by using Dict Gym space. For data that varies with time (stringency, normalized GDP, $R_{e}$) we use a simple long short-term memory architecture [60]. For other data, such as the current proportion of the population that is susceptible, infected, and recovered, we use a simple fully connected layer. The output from both of these networks is concatenated and used by the reinforcement learning agent for training. We train the model for $2742$ time steps and some of the best results are presented.
@@ 3 Results
Using the simple SIR model from Equations 114, to model the disease dynamics we get fig. 4. Here, it can be observed that the SIR model accurately fits the susceptible population and recovered population but overestimates the infected population by a significant margin which can create complications. This is because disease dynamics are controlled by this population and our work involves rewarding the agent when the proportion of infected individuals falls below a predetermined threshold. Therefore, an overestimation of the infected population could lead to incorrect decision-making and undesirable outcomes.
@@
Figure 4: SIR Model Comparison for India. The figure presents a comparison between the fitted simple SIR model (Equations 114) and real data. Here an evident overestimation of the infected population is observed.
Combining the lockdown dynamics in the SIR model using Equations 1522, we get the following fig. 5. Here, it can be observed there’s an overestimation of infected individuals, but, the two stages of the epidemic are being accounted for. This is what suggests that there might be depletion of infected individuals through vaccination.
@@
Figure 5: SIR Model with Lockdown Analysis for India. This figure illustrates the fitting of the SIR model with lockdown (Equations 1522) in comparison to real data. The introduction of lockdown measures showcases discernible effects on the dynamics of disease progression. While an overestimation persists, the model’s peaks now closely align with the observed data and are able to capture key trends.
Incorporating vaccination dynamics into the SIR model with lockdown measures, as described by Equations 2331, we get the following fig. 6. Here, because the value of $\nu$ eq. 28 is negligible, it doesn’t change the results significantly compared to the previous model (Equations 1522 and fig. 5). Therefore, a time-varying $\nu$ shall be able to better account for the these dynamics.
@@
Figure 6: SIR Model with Lockdown and Vaccination for India. This figure displays the fitting of the SIR model with lockdown (Equations 2331) compared to the real data. The infection trends closely resemble those depicted by the SIR model with lockdown, as illustrated in fig. 5(c) and this is because the rate of vaccination is negligible eq. 28. This is suggestive of a rate of vaccination that varies with time.
For SIR model with lockdown and time-varying vaccination rate from Equations 3543, we get the following fig. 7. With a time-varying $\nu$ (vaccination rate) and the effect of lockdown, our model is able to account for the infected individuals and reduce the cost in comparison to all the previously formalized models for the data. This shows how interventions and changes in the way people behave in response of an epidemic [13] play a major role in the way the epidemic unfolds.
@@
Figure 7: SIR Model with Lockdown and Time-varying Vaccination Rate. This figure displays the fitting of the SIR model with lockdown and time-varying vaccination rate (Equations 3543) compared to the real data. Incorporating a time-varying vaccination rate enhances the model’s ability to capture variations in the infected population over time.
@@
While non-pharmaceutical interventions (NPIs) can effectively manage the epidemic, they impose economic burdens on developing nations. In fig. 9, we plot the normalized GDP against the stringency and calculate various metrics like the Pearson correlation coefficient, coefficient of determination ($r^{2}$), and p-value for three countries (India, Mexico, Brazil) which are Emerging Market and Developing Economies [4] from May 2020 to October 2022. It can be observed from fig. 9 that strict policies have a negative effect on the normalized GDP in these economies. However, this trend is not uniformly seen in advanced economies like the USA, Japan, and Canada as shown in fig. 10. In these countries, other factors could be contributing to the decrease in normalized GDP besides the implementation of stricter policies.
@@
Figure 10: Stringency and GDP for Advanced Economies. Here, “(actual)” is the real data, “(modelled)” is the model for normalized GDP given the stringency. In economically advanced nations, when modeling the normalized GDP against stringency measures, we observe substantial p-values, providing evidence against the null hypothesis. However, the significance levels are not as high as those found for developing economies (Figures 9(a)9(f)). This is reflected in the lower R-squared scores, which indicates that the relationship between these variables may be less pronounced when compared to developing economies.
@@
After median filtering to smooth the output (to reinforce the negative reward from changing the stringencies) from the trained reinforcement learning agent, here are some of the results obtained. In the presented result fig. 11, we can see the reinforcement learning agent outperform the modelled outcome. A strategic decision is made by the agent to maintain the stringency index below 80 after the April, 2020 fig. 11(a). This approach allows for the natural progression of disease dynamics, resulting in a rapid reduction of the effective reproduction number $R_{e}$ to below $1.2$ after October, 2020 (refer to figure fig. 11(e)). After October, 2021 there’s a decrease in the stringency which leads to an increase in the normalized GDP, indicating an economic upturn. While this strategy poses a higher risk in terms of infection rates during the initial phase of the epidemic (prior to vaccine rollout) as well as the later phase (second peak of infected individuals fig. 11(c)), but it proves to be more beneficial for the nation’s economy in the long run. Despite the economic benefits in the long run, this strategy is not the most effective for the government to adopt due to the high number of infected individuals. Therefore, we propose an alternative strategy that involves some loss in terms of the economic impact.

### v2

@@ 1 Introduction
The need for a nuanced understanding of how interventions impact both health outcomes and economic indicators became increasingly evident, prompting a comprehensive examination by epidemiologists to assist policymakers [7]. The outbreak of COVID-19 has prompted epidemiologists to research on various aspects, including mobility control [8, 9], vaccination strategies [10, 11], non-pharmaceutical interventions (NPIs) like restricting population movements and gatherings, closing schools and businesses, requiring masks indoors [12, 13, 14], and financial considerations [15]. Despite the numerous studies conducted, very few explore how common interventions meet multiple policy objectives or how a precise articulation of the main policy goals directs the selection of the most effective interventions in terms of health and economic results [16, 8, 17, 18, 19, 20, 21, 22]. The economic impact of the COVID-19 pandemic varied between rich and poor countries. Although COVID-19 deaths had a slightly larger negative effect on the Gross Domestic Product (GDP) in advanced economies, this difference was not statistically significant. However, lockdown restrictions were found to have a more damaging impact on economic activity in emerging and developing economies [6, 23, 24]. It’s also suggested that an increase in COVID-19 cases was associated with the introduction of harsher NPIs and lockdown measures could be relaxed once vaccination rates increase [23, 25].
@@
Many economists have studied the effect of COVID-19 on the economy of nations [6, 26, 27, 28]. In advanced economies like Korea, where the stringency index was below the median, the recession was milder than other advanced economies like the United Kingdom where the stringency was much higher [26], they achieved it mostly with very aggressive testing, contact tracing, and enforced quarantines [29, 30]. In India, social distancing and containment measures have been effective in reducing the number of COVID-19 cases but have come with economic costs. Social distancing had the most adverse effect on the economy in areas with high urbanization [27].
@@
In this paper, we optimize the government policies regarding stringency as it controls both the spread of the disease and the economy. To model the epidemiological data [31] we use the simple SIR model without vital dynamics [32, 33, 34], as it is assumed that the timescale is small enough that it can be neglected [35]. By lesioning the model, as opposed to proposing a new mathematical model with more specialized compartments to more accurately represent the actual environment [36, 37], we effectively model the disease progression. Our model (SIR with lockdown and time-varying vaccination rate) builds on the foundational SIR model, by accounting for the recovery reached through vaccination [38, 39, 40, 41, 42] and the effects of lockdown [43, 44, 21, 45]. Although the traditional SIR model is a valuable tool for understanding the spread of infectious diseases, it assumes that parameters like the transmission rate ($\beta$) and recovery rate ($\gamma$) are constant over time, which may not always be the case. In this paper, we propose a more sophisticated approach by introducing a time-dependent SIR model [46], enabling us to account for the changing dynamics of the pandemic due to factors such as lockdowns and vaccination rates. This proposed model effectively addresses the real-world conditions acts as a solution that is both effective and extendable. However, the study has limitations; First, the deterministic SIR model (predecessor to our proposed model) fails to account for chance in disease spread and lacks confidence intervals on results and while stochastic models incorporate chance, they are typically more challenging to analyze than their deterministic counterparts [33]; Secondly, the underreporting of cases during the period selected by our study; Lastly, the reinforcement learning agent should be resistant to how the vaccination rate changes and different values for $\beta$ and $\gamma$ – keeping them the same scopes the environment for succumbing to wishful thinking which can be potentially dangerous. Therefore, before an actual deployment of the model, it would be a good measure to introduce stochasticity to these parameters ($\beta$ and $\gamma$) and the vaccination rate ($\nu$).
@@
In summary, by using reinforcement learning augmented with deep learning techniques for the SIR with lockdown and time-varying vaccination rate environment, we can better understand the effects of lockdown measures on both public health outcomes and economic stability during a pandemic. However, it is crucial to consider the limitations of this approach and take into account a comprehensive set of decision factors in order to make informed policy decisions that are tailored to specific socio-economic contexts.
@@ 2.1 Simple SIR Model
During an epidemic, the fundamental drivers of an epidemic growth is the rate of infection $\beta$, i.e., the average number of infections per infected case and the infectious period $1/\gamma$, i.e., the average period for which the infected case is infected for. Epidemics can only happen if the case is infectious enough for long enough and this defined by $R_{0}=\beta/\gamma$. Here, $R_{0}$ is the average number of secondary infections caused by each infected case, in an otherwise fully susceptible population.
At the peak of an epidemic, there is a decline as there are no more susceptible people left in the pool, therefore, $R_{e}$ (effective reproductive number) comes into play. $R_{e}$ is defined as the average number of secondary cases arising from an infected case, at a given point in an epidemic, therefore, it takes into account the existing immunity of the system [62].
@@
$S$ is the number of susceptible people and $N$ is the total population. At the start of an epidemic when everyone is susceptible, $R_{e}=R_{0}$ as, $S=N$ (i.e., the whole population is susceptible). $\beta$ and $\gamma$ are also used to define probability of and infectious individual infecting another individual $\beta/(\beta+\gamma)$ and the probability of recovery, $\gamma/(\beta+\gamma)$.
Most government policies look at the value of $R_{e}$ to come up with an effective strategy to combat the disease as the fate of the evolution of the disease depends upon it. When $R_{e}$ is less than one, the infected population $I$ will steadily decline to zero. Conversely, if $R_{e}$ is greater than one, the infected population will increase. In other words, when $\frac{dI(t)}{dt}<0\Rightarrow R_{e}<1$ and $\frac{dI(t)}{dt}>0\Rightarrow R_{e}>1$, therefore, the effective reproductive rate $R_{e}$ serves as a critical threshold that determines whether an infectious disease will rapidly extinguish or escalate into an epidemic [35].
To estimate the parameters $\beta$ and $\gamma$ for India based on the data from May 2020 to October 2022, we simply define two cost functions eqs. 8 and 9 to calibrate the model with the use of Huber loss [63]. Our model typically uses eq. 8, considering all three compartments: susceptible, infected and recovered. However, there are instances where we need to balance modeling all groups and focusing on the infected group (population that drives disease spread). In such cases, we consider losses from both eqs. 8 and 9. A weighted sum of the loss functions allows for trade-offs between comprehensive modeling and focusing on infected group dynamics (see section 2.4, where a hyperparameter (window length) is selected keeping both these losses in mind).
@@
Where, $S$ is the number of susceptible people and $\hat{S}$ is the predicted number of susceptible people, similarly, $I$ for infected and $\hat{I}$ for the predicted number of infected people, and $R$ for recovered. Using the equations eqs. 13, we minimize the cost function eq. 13 using the Nelder-Mead method [64] to estimate the parameters like $\beta$, $\gamma$ and fit the model to actual data. The following parameters and loss is obtained:
@@ 2.2 SIR Model with Lockdown
Now that a simple SIR model has been established – we need to model the effects of the stringency index (measure for the strictness of lockdown) on $\beta$ (the effective transmission rate). To do this, we say the flow of susceptibles not only depend on $\beta$ but also $s(t)$ the stringency index at time [45, 65, 44, 66, 22]. The stringency index is a composite measure based on nine response indicators including school closures, workplace closures, and travel bans, rescaled to a value from 0 to 100 (100 = strictest) [67]. This index simply records the strictness of government policies and does not measure or imply the appropriateness or effectiveness of a country’s response, i.e., a higher score does not necessarily mean that a country’s response is “better” than others lower on the index.
@@ 2.4 Optimizing Window Length for Time-varying Vaccination Rate
However, as observed by the value of $\nu_{optimal}$ from eq. 28, which is almost negligible and the overestimation of infected individuals in fig. 6(b) suggests that $\nu$ might be varying with time. This suggests to accurately estimate the infected population a time-varying vaccination rate should be used as the transition from susceptibility to direct recovery fluctuates with time [24, 38]. We first find the optimal window length [68] for which the value of $\nu$ is constant for a time sub-interval and results in the least loss from eqs. 8 and 9. For this we use different window lengths ($window\_lengths=5,10,15\dots 40,45,50\textrm{ days}$).
@@
Using the model described by eqs. 3234 for each $window\_length_{i}$ in $window\_lengths$, where $i=1,2,3\dots 10$, we calculate the $time\_varying\_\nu$. For each $time\_varying\_\nu$ we then calculate the $loss\_SIR$ and $loss\_I$. This is done as follows:
@@
Figure 1: Loss for Different Window Lengths. We try different window lengths to find the optimal loss for both cases, either when predicting all three populations (susceptibles, infected, recovered) or just the infected population.
The results in fig. 1 indicate that a window length of $10$ days yields the least overall loss for all three population groups. However, this window length results in a poor approximation for just the infected group, which is crucial for accurately modelling the spread of the disease. Consequently, we have decided to use a window length of $15$ days, which provides a more accurate approximation for the infected population while still maintaining reasonable loss for the other groups.
@@
Figure 2 shows us that the $\nu$ coincides with the actual of data of when the vaccination drive was first launched in India [69]. Therefore, using these values we finally recompute $\beta_{optimal}$ and $\gamma_{optimal}$ by supplying them into the equations for the SIR Model with lockdown and time-varying $\nu$.
@@ 2.6 Modelling Normalized GDP with Stringency
Now that a relation between $\beta$ and $s(t)$ is set up, it must be investigated how stringency index affects the normalized GDP [70, 71]. To do this, a polynomial equation of the third degree is fitted to the data points $f(x)=ax^{3}+bx^{2}+cx+d=y$, here, $x$ is the stringency ($s$) and $y$ the normalized GDP, and we minimize the squared error to find the values of coefficients $a,b,c,d$. For India after fitting a third-degree polynomial, the following equation is obtained:
@@ 2.7 Reinforcement Learning
  • Actions $\altmathcal{A}$: The stringency index variable was analyzed with a sample size of 915. The mean value was approximately $61.96505$, with a standard deviation of $17.66983$. The minimum value was $31.48$, while the maximum value reached $96.3$. And the differences between two consecutive stringencies had a mean of $-0.070919$, and standard deviation of $1.42715$, with the minimum being $-14.36$ and maximum $16.67$. Based on this, we define the discrete action space. There are 7 actions for the agent, it can keep the stringency index same, reduce/increase by 2.5, reduce/increase by 5, and reduce/increase by 10 given that the stringency index doesn’t exceed 100 or go below 0.
@@ 2.7.1 Defining the Reward Function
The stringency index emerges as a critical factor influencing both the normalized GDP and the rate of infection spread. The decision to escalate or de-escalate the stringency index is a strategic one, with significant implications. Increasing the stringency decreases the spread of the infection. Conversely, it must be noted that herd immunity can only be achieved when the epidemic reaches its peak, i.e., when the effective reproductive number is equal to one ($R_{e}=1$). This can only happen by lowering the stringency index which would allow the natural dynamics of the epidemic to transpire such that the population of susceptible individuals has depleted enough such that it is insufficient to propagate the disease further. Therefore, stringency is used to control the number of infected people and slow down the rate at which the epidemic reaches its peak, so that hospitals could house the number of infected people.
In reinforcement learning, positive rewards promote and negative rewards demote actions. The agent tries to generate such a policy/knowledge to avoid the discouraging situation by following the policy. By designing a proper reward function, it is possible to generate such an agent that may follow the human desired situation. While designing a reward function, it is important to note that the rewards we set up truly indicate what we want accomplished. In particular, the reward signal is not the place to impart to the agent prior knowledge about how to achieve what we want it to do [52]. Taking inspiration from similar work [19], we define the reward function.
@@ 2.7.2 Deep Reinforcement Learning and Training
The agent observes the percentage of the population that is susceptible, infected, recovered, and time-varying data like the GDP, and previous actions taken to change the stringency, and $R_{e}$. Since Stable Baselines3 can support multiple inputs (time-series data, single data points and images) by using Dict Gym space. For data that varies with time (stringency, normalized GDP, $R_{e}$) we use a simple long short-term memory architecture [60]. For other data, such as the current proportion of the population that is susceptible, infected, and recovered, we use a simple fully connected layer. The output from both of these networks is concatenated and used by the reinforcement learning agent for training. The schematic diagram of neural networks used to inform reinforcement learning agent’s decision-making process is given in fig. 3. We train the model for $2742$ time steps and some of the best results are presented.
@@ 3 Results
Using the simple SIR model from eqs. 114, to model the disease dynamics we get fig. 4. Here, it can be observed that the SIR model accurately fits the susceptible population and recovered population but overestimates the infected population by a significant margin which can create complications. This is because disease dynamics are controlled by this population and our work involves rewarding the agent when the proportion of infected individuals falls below a predetermined threshold. Therefore, an overestimation of the infected population could lead to incorrect decision-making and undesirable outcomes.
@@
Figure 4: SIR Model Comparison for India. The figure presents a comparison between the fitted simple SIR model (eqs. 114) and real data. Here, an evident overestimation of the infected population is observed.
Combining the lockdown dynamics in the SIR model using eqs. 1522, we get the following fig. 5. Here, it can be observed there’s an overestimation of infected individuals, but, the two stages of the epidemic are being accounted for. This is what suggests that there might be depletion of infected individuals through vaccination.
@@
Figure 5: SIR Model with Lockdown Analysis for India. This figure illustrates the fitting of the SIR model with lockdown (eqs. 1522) in comparison to real data. The introduction of lockdown measures showcases discernible effects on the dynamics of disease progression. While an overestimation persists, the model’s peaks now closely align with the observed data and is able to capture key trends.
Incorporating vaccination dynamics into the SIR model with lockdown measures, as described by eqs. 2331, we get the following fig. 6. Here, because the value of $\nu$ eq. 28 is negligible, it doesn’t change the results significantly compared to the previous model (eqs. 1522 and fig. 5). Therefore, a time-varying $\nu$ shall be able to better account for the these dynamics.
@@
Figure 6: SIR Model with Lockdown and Vaccination for India. This figure displays the fitting of the SIR model with lockdown (eqs. 2331) compared to the real data. The infection trends closely resemble those depicted by the SIR model with lockdown, as illustrated in fig. 5(c) and this is because the rate of vaccination is negligible eq. 28. This is suggestive of a rate of vaccination that varies with time.
For SIR model with lockdown and time-varying vaccination rate from eqs. 3543, we get the following fig. 7. With a time-varying $\nu$ (vaccination rate) and the effect of lockdown, our model is able to account for the infected individuals and reduce the cost in comparison to all the previously formalized models for the data. This shows how interventions and changes in the way people behave in response of an epidemic [13] play a major role in the way the epidemic unfolds.
@@
Figure 7: SIR Model with Lockdown and Time-varying Vaccination Rate. This figure displays the fitting of the SIR model with lockdown and time-varying vaccination rate (eqs. 3543) compared to the real data. Incorporating a time-varying vaccination rate enhances the model’s ability to capture variations in the infected population over time.
@@
While non-pharmaceutical interventions (NPIs) can effectively manage the epidemic, they impose economic burdens on developing nations. In fig. 9, we plot the normalized GDP against the stringency and calculate various metrics like the Pearson correlation coefficient, coefficient of determination ($r^{2}$), and p-value for three countries (India, Mexico, Brazil), which are Emerging Market and Developing Economies [4] from May 2020 to October 2022. It can be observed from fig. 9 that strict policies have a negative effect on the normalized GDP in these economies. However, this trend is not uniformly seen in advanced economies like the USA, Japan and Canada as shown in fig. 10. In these countries, other factors could be contributing to the decrease in normalized GDP besides the implementation of stricter policies.
@@
Figure 10: Stringency and GDP for Advanced Economies. Here, “(actual)” is the real data, “(modelled)” is the model for normalized GDP given the stringency. In economically advanced nations, when modeling the normalized GDP against stringency measures, we observe substantial p-values, providing evidence against the null hypothesis. However, the significance levels are not as high as those found for developing economies (figs. 9(a)9(f)). This is reflected in the lower R-squared scores, which indicates that the relationship between these variables may be less pronounced when compared to developing economies.
@@
After median filtering to smooth the output (to reinforce the negative reward from changing the stringencies) from the trained reinforcement learning agent, here are some of the results obtained. In the presented result fig. 11, we can see the reinforcement learning agent outperform the modelled outcome. A strategic decision is made by the agent to maintain the stringency index below 80 after April, 2020 fig. 11(a). This approach allows for the natural progression of disease dynamics, resulting in a rapid reduction of the effective reproduction number $R_{e}$ to below $1.2$ after October, 2020 (refer to figure fig. 11(e)). After October, 2021 there’s a decrease in the stringency which leads to an increase in the normalized GDP, indicating an economic upturn. While this strategy poses a higher risk in terms of infection rates during the initial phase of the epidemic (prior to vaccine rollout) as well as the later phase (second peak of infected individuals fig. 11(c)), but it proves to be more beneficial for the nation’s economy in the long run. Despite the economic benefits in the long run, this strategy is not the most effective for the government to adopt due to the high number of infected individuals. Therefore, we propose an alternative strategy that involves some loss in terms of the economic impact.